OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (1, 1) and has a y-intercept of 2.

OpenStudy (anonymous):

 It's simply \[y=-x\]

OpenStudy (haseeb96):

m=y2-y1\x2-x1 (x1,y1)=(1,1) (x2,y2)=(0,2) m=-1 y=mx+b so but m in it y=-x+b this is the general equation

OpenStudy (haseeb96):

i hope that you can understand by this method

OpenStudy (anonymous):

I do, thankyou so much(: